OpenStudy (anonymous):

Find the length of the curve. Note: You will need to evaluate your integral numerically. Round your answer to one decimal place. x = cos(4t), y = sin(3t) for 0 ≤ t ≤ 2π Please help/explain this vector calc problem!

OpenStudy (anonymous):

In general, the formula for the length of a formula which is specified through \(x=x(t)\) and \(y=y(t)\) in the interval \(a\lt t\lt b\) is: \[L=\int^b_a\sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}\] So in this case we have: \[L=\int^{2\pi}_0\sqrt{1+\frac{1}{16}cos^2(4t)+\frac{1}{9}sin^2(3t)}\]

OpenStudy (tkhunny):

?? \(\dfrac{d}{dt}\cos(4t) = -4\sin(4t)\) \(\dfrac{d}{dt}\sin(3t) = 3\cos(3t)\)

OpenStudy (anonymous):

Thank you both so much! :)